# Edinburg, Texas, USA

### v1

@@ Abstract
Neuronal oscillations are related to symptoms of Parkinson’s disease. The random inputs could affect such oscillations in the brain states that translate collective activities of neurons interconnected via synaptic connections. In this paper, we study coupled effects of channels and synaptic dynamics under the stochastic influence, together with spike-timing-dependent plasticity (STDP) of healthy brain cells with applications to Parkinson’s disease (PD). In particular, we investigate the effects of random inputs and input correlations in a subthalamic nucleus (STN) cell membrane potential model. Our numerical results show that the random inputs strongly affect the spiking activities of the STN neuron not only in the case of healthy cells but also in the case of PD cells in the presence of DBS treatment. The STDP increases the interspike interval (ISI) regularity of spike trains of the output neurons. However, the existence of a random refractory period and random input current in the system may substantially influence an increased irregularity of spike trains of the output neurons. Furthermore, the presence of the stochastic influence together with spike-timing-dependent plasticity could increase the correlation of the neurons. These effects would potentially contribute to the management of PD symptoms.
@@ 1 Introduction
One of the most common neurodegenerative disorders of aging is Parkinson’s disease (PD), which is a progressive neurological degenerative disorder characterized by a severe loss of dopaminergic neurons in the substantia nigra pars compacta. PD is characterized by cardinal motor symptoms including slowness of movement (bradykinesia), akinesia, rigidity, and tremors at rest. Neuronal oscillations and symptoms of Parkinson’s disease (PD) are closely connected [20, 7, 34]. Many different treatments focus on the subthalamic nucleus (STN) to improve such motor symptoms, for instance, ablation surgery of STN or its fiber connections. Deep brain stimulation (DBS) of the STN has recently become an effective therapy for PD [20]. The DBS is a very impressive method [35] due to the fact that PD, characterized by the inadequacy of a chemical substance in the brain, can also be successfully treated with the passage of only electrical currents without a concomitant supply of biological or chemical reactions/factors. In general, biological neurons in the brains transmit information by generating spikes and such neurons are connected by synapses that process and store information. To better understand brain activities, we need to know how synapses work. In general, chemical synapses are more common in the brain, and these synapses do not physically join neurons [44, 39]. To get closer to the real scenarios of brain activities, one needs to account for models of a synapse in which its synaptic strength changes as a function of the relative timing (i.e., time difference) between the spikes of the presynaptic and postsynaptic neurons, respectively. This change in the synaptic weight is known as spike-timing-dependent plasticity (STDP).
Many models have been proposed to analyze the dynamics of synaptic coupling of human brains in neurodegenerative disorders and therapeutic targets for such diseases. A model of STDP mediated by dopamine and its role in PD pathophysiology has been reported in [30]. For the first time, the results provided in [20] have demonstrated that neuronal oscillations are predictive of DBS outcome. The DBS therapy for the noisy cortex and irresponsive subthalamus to improve parkinsonian locomotor activities has been reported in [26]. The authors in [43] have investigated the coupled effects of channels and synaptic dynamics in stochastic modelling of healthy and Parkinson’s-disease-affected brains. In [48], the authors have found that eSTDP and iSTDP restrict the balance of the excitatory–inhibitory balanced neuronal network and play a fundamental role in maintaining network stability and synchronization. Furthermore, their results would potentially also contribute to the guidance of the treatments and further research of neurodegenerative diseases. Moreover, the results reported in [36] show that STDP can support the multiplexing of rhythmic information. Additionally, STDP could demonstrate how to retain functionality despite continuous remodeling of all the synaptic weights.
The computational aspects of sensory information processing span from individual cells to the network level. The physical properties of single and small networks of nerve cells enable them to process and store information. Understanding of the cellular and dendritic mechanisms that could contribute to the processing of sensory information in single neurons has also been greatly increased. However, still very little is known about how the biophysical properties of single neurons are actually used to implement specific computations. Two types of neuronal computations thought to be fundamental to processing information within the nervous system are the multiplication of independent signals and invariance of neuronal responses. The evidence of a behaviorally induced form of synaptic plasticity that would encourage the development of fine-scale structured input patterns and the binding of features within single neurons has been presented in [31]. Feedforward inhibition is not only traditionally thought to sharpen the responses and temporal tuning of feedforward excitation onto principal neurons but also could also convey information used by single neurons to implement dendritic computations on sensory stimulus variables [47]. A discussion on the discrimination of the broad spatial statistics of synaptic inputs within the dendrites of a single neuron has been reported in [11]. Using the description based on single and dual dendritic recordings in vivo together with computational modelling, the membrane impedance of a collision detection neuron in the grasshopper Schistocerca americana has been characterized in [10]. A model of active dendritic integration and mixed neocortical network representations during an adaptive sensing behavior for single neurons has been considered in [32].
In the real scenarios in the neuronal models, the random fluctuations could affect the spiking activities of the neurons. In particular, stochastic factors arise through sensory fluctuations, brainstem discharges and thermal energy or random fluctuations at a microscopic level, such as the Brownian motion of ions. Such random fluctuations could cause burst discharges in brain systems. The brain rhythm bursts are enhanced by multiplicative noise [2]. The authors in [40] have investigated the effects of noise in gamma oscillations in a model of neuronal networks with different reversal potentials. Specifically, the presence of noise in a neuronal system can also be a solution in information processing [1, 44]. The authors in [17] have shown that the detectability of weak signals in nonlinear systems (known as stochastic resonance) can be enhanced by random noise. Furthermore, they have also pointed out that the phase-based simplification of the STN neurons can accurately predict responses to temporally complex trains of inputs even when the perturbations in timing are large enough to obscure the oscillatory nature of the neuron’s firing. The question ”Is this variation an unavoidable effect of generating spikes by sensory or synaptic processes (’neural noise’) or is it an important part of the ’signal’ that is transmitted to other neurons?” has been discussed in [38]. In general, the occurrence of synchronous events can emerge from the variety of mechanisms in physical systems. In neuronal networks, signal-evoked and intrinsic noise spike correlations both originate in the intricate connectivity of a neuronal network. Each cortical neuron receives inputs from approximately $10^{4}$ other neurons and sends out signals via its synapses to about $10^{4}$ others. In such a neuronal network, spiking cross-correlations in the activity of two neurons can emerge from direct synaptic connections or shared presynaptic partners. Because neurons are highly interconnected, it is almost unavoidable that two neurons in a network share some of their inputs. Just as multiple synaptic interaction structures can potentially give rise to the same pairwise correlations, a particular functional form of the spike correlation function may not be a unique signature of interneuronal interactions. In [21], the authors have shown that pairs of neurons receiving correlated input also exhibit correlations arising from precise spike-time synchronization by using simulations and experiments in rat hippocampal neurons. The role of input correlations in shaping the variability and noise correlations of evoked activity in the neocortex has been investigated in [3]. The effect of interpopulation STDP on synchronized rhythms in neuronal networks with inhibitory and excitatory populations has been discussed in [24]. The authors in [48] have consider the synchronization and oscillation behaviors of excitatory and inhibitory populations with STDP.
This paper is built upon and extend the recent works [44, 39, 43], where the authors considered the static synapses whose weights are fixed and the dynamic synapses whose synaptic strength is dependent on the recent spike history (short-term plasticity). However, in real brain models, the strength of connections between neurons in the brain is based on the relative timing of a particular neuron’s output and input action potentials. Drawing from the fields of PD studies and the effects of natural random factors in biological system dynamics, we develop and investigate a model of coupled effects of channels and synaptic dynamics by using stochastic modelling of healthy brain cells with applications in PD. In particular, we consider a cell membrane potential model in the STN part of the human brain. Our analysis focuses on considering a Langevin stochastic equation in a numerical setting for a cell membrane potential with random inputs and input correlations under the influence of STDP. We discuss the role of STDP in neuromorphic systems. Next, we provide numerical examples and discuss the effects of random inputs on the time evolution of the STN cell membrane potential as well as the spiking activities of the STN neuron. Our numerical results indicate that STDP enhances the regularity of interspike intervals (ISI) in the spike trains of output neurons. However, the presence of a random refractory period, along with random input currents, may significantly increase the irregularity of these spike trains. Additionally, the combination of stochastic influences and STDP could enhance the correlation among neurons. These effects may contribute to the management of Parkinson’s disease symptoms.
@@ 2 Model description
The phenomenology of STDP is generally described as a biphasic exponentially decaying function. That is, the instantaneous change in weights is given by:
@@
where $\Delta W$ denotes the change in the synaptic weight, $A_{+}$, $A_{-}$ determine the maximum amount of synaptic modification (which occurs when the timing difference between presynaptic and postsynaptic spikes is close to zero), while $\tau_{+}$ and $\tau_{-}$ represent the ranges of pre-to-postsynaptic interspike intervals over which synaptic strengthening or weakening occurs. Thus, $\Delta W>0$ means that the postsynaptic neuron spikes after the presynaptic neuron.
This model captures the phenomena that repeated occurrences of presynaptic spikes within a few milliseconds before postsynaptic action potentials lead to long-term potentiation (LTP) of the synapse, whereas repeated occurrences of presynaptic spikes after the postsynaptic ones lead to long-term depression (LTD) of the same synapse.
The latency between presynaptic and postsynaptic spike ($\Delta t$) is defined as:
@@
where $t_{\text{pre}}$ and $t_{\text{post}}$ are the timings of the presynaptic and postsynaptic spikes, respectively.
@@ 2.1 Keeping track of pre- and postsynaptic spikes
Since a neuron will receive numerous presynaptic spike inputs, in order to implement STDP by taking into account different synapses, we first have to keep track of the pre- and postsynaptic spike times throughout the simulation.
A convenient way to do this is to define the following equation for each postsynaptic neuron:
$\tau_{-}\frac{dM}{dt}=-M$ (4)
and whenever the postsynaptic neuron spikes, we have
@@
This way $M(t)$ tracks the number of postsynaptic spikes over the timescale $\tau_{-}$.
Similarly, for each presynaptic neuron, we define:
$\tau_{+}\frac{dP}{dt}=-P$ (6)
and whenever there is a spike on the presynaptic neuron, we have
@@
The variables $M(t)$ and $P(t)$ are very similar to the equations for the synaptic conductances, i.e., $g_{i}(t)$, except that they are used to keep track of pre- and postsynaptic spike times on a much longer timescale. Note that, $M(t)$ is always negative, and $P(t)$ is always positive. You can probably already guess that $M$ is used to induce LTD and $P$ to induce LTP because they are updated by $A_{-}$ and $A_{+}$, respectively.
Important note: $P(t)$ depends on the presynaptic spike times. If we know the presynaptic spike times, $P$ can be generated before simulating the postsynaptic neuron and the corresponding STDP weights.
@@ 2.2 Implementation of STDP
We will change the value of the peak synaptic conductance based on the presynaptic and postsynaptic timing, thus using the variables $M(t)$ and $P(t)$.
Each synapse $i$ has its own peak synaptic conductance ($\bar{g}_{i}$), which may vary between $[0,\bar{g}_{\text{max}}]$, and will be modified depending on the presynaptic and postsynaptic timing.
@@
    When the $i$th presynaptic neuron elicits a spike, its corresponding peak conductance is updated according to the following equation:
@@
    Note that, $M(t)$ tracks the time since the last postsynaptic potential and is always negative. Hence, if the postsynaptic neuron spikes shortly before the presynaptic neuron, the above equation shows that the peak conductance will decrease.
@@
    When the postsynaptic neuron spikes, the peak conductance of each synapse is updated according to:
@@
    Note that, $P(t)$ tracks the time since the last spike of $i$th pre-synaptic neuron and is always positive.
    Thus, the equation given above shows that if the presynaptic neuron spikes before the postsynaptic neuron, its peak conductance will increase.
@@ 2.3 The leaky integrate-and-fire neuron connected with synapses that show STDP
We connect $N$ presynaptic neurons to a single postsynaptic neuron. We do not need to simulate the dynamics of each presynaptic neuron as we are only concerned about their spike times. So, we will generate $N$ Poisson-type spikes. Here, we will assume that all these inputs are excitatory.
@@
where $I_{\text{app}}$ is the external input current, while $C_{m}$ is the membrane capacitance and $t\in[0,T]$. Additionally, in (13), $V_{\text{th}}$ denotes the membrane potential threshold to fire an action potential. In this model, we assume that a spike takes place whenever $V(t)$ crosses $V_{\text{th}}$ in the STN membrane potential. In that case, a spike is recorded and $V(t)$ resets to $V_{\text{reset}}$ value. Hence, the reset condition is summarized by $V(t)=V_{\text{reset}}$ if $V(t)\geq V_{\text{th}}$. The quantity $I_{\text{ahp}}$ represents the calcium-activated potassium current for the spike after hyperpolarization in STN.
@@
In our model (17)–(29), as we mentioned above, we use the simplest input spike train with Poisson process in which the stochastic process of interest provides a suitable approximation to stochastic neuronal firings [42]. The input spikes will be carried out by the quantity $\sum_{k}\delta(t-t_{k})$ in the equation (10) and the input spikes are given when every input spike arrives independently of other spikes. The process will be described as follows:
@@
In general, the information on stimulating activities in a neuron can be provided by the irregularity of spike trains. The time interval between adjacent spikes is called the ISI. The coefficient of variation (CV) of the ISI in a cell membrane potential with multiple inputs can bring useful information about the output of a decoded neuron. In what follows, we will demonstrate that when we increase the value of $\sigma_{\text{ref}}$, the irregularity of the spike trains increases (see also [12]). The spike irregularity of spike trains can be described via the coefficient of variation of the inter-spike-interval (see, e.g., [5, 12]) as follows:
@@
Current Gating variables Gating variables Parameters $I_{\text{L}}=g_{L}(v-E_{\text{L}})$ $g_{L}=2.25$ (nS) $E_{\text{L}}=-60$ (mV) $I_{\text{Na}}=g_{\text{Na}}m_{\infty}^{3}(V)h(V)(V-E_{\text{Na}})$ $m_{\infty}(V)=1/(1+\exp(-\frac{V+30}{15}))$ $h_{\infty}(V)=1/(1+\exp(-\frac{V+39}{3.1})$ $g_{\text{Na}}=37$ $\tau_{h}(V)=1+500/(1+\exp(-\frac{V+57}{-3})$ $E_{\text{Na}}=55$ (mV) $I_{\text{K}}=g_{\text{K}}n^{4}(V)(V-E_{\text{K}})$ $n_{\infty}(V)=1/(1+\exp(-\frac{V+32}{8}))$ $g_{\text{K}}=45$ (nS) $\tau_{n}(V)=1+100/(1+\exp(-\frac{V+80}{-26})$ $E_{\text{K}}=-80$ (mV) $I_{\text{T}}=g_{\text{T}}a^{3}_{\infty}(V)b_{\infty}^{2}(r)r(V)(V-E_{\text{T}})$ $a_{\infty}(V)=1/(1+\exp(-\frac{V+63}{7.8}))$ $r_{\infty}(V)=1/(1+\exp(\frac{V+67}{2})$ $g_{\text{T}}=0.5$ (nS) $b_{\infty}(V)=1/(1+\exp(-\frac{V-0.4}{0.1}))$ $\tau_{r}(V)=7.1+17.5/(1+\exp(-\frac{V+68}{-2.2})$ $E_{\text{T}}=0$ (mV) $-1/(1+\exp(4))$ $I_{\text{Ca}}=g_{\text{Ca}}c^{2}(V)(V-E_{\text{Ca}})$ $c_{\infty}(V)=1/(1+\exp(-\frac{V+20}{8}))$ $g_{\text{Ca}}=2$ (nS) $\tau_{c}(V)=1+10/(1+\exp(\frac{V+80}{26})$ $E_{\text{Ca}}=140$ (mV) $I_{\text{ahp}}=g_{\text{ahp}}(V-E_{\text{ahp}})(\frac{\text{Ca}}{\text{Ca}+15})$ $g_{\text{ahp}}=20$ (nS) $E_{\text{ahp}}=-80$ (mV) $I_{\text{dbs}}=5+5\sin(2\pi t)$ (pA)

### v2

@@ Abstract
Neuronal oscillations are closely related to the symptoms of Parkinson’s disease (PD). In this study, we explore how random fluctuations (or "stochastic inputs") affect these oscillations in brain states, which reflect the collective activity of interconnected neurons. These random inputs are modeled in the context of the subthalamic nucleus (STN), a brain region implicated in PD, and their interaction with synaptic dynamics and spike-timing-dependent plasticity (STDP) in both healthy and PD-affected neurons. Specifically, we investigate the effects of random synaptic inputs and their correlations on the membrane potential of STN neurons. Our results show that these random inputs significantly influence the firing patterns of STN neurons, both in healthy cells and in those affected by PD under deep brain stimulation (DBS) treatment. We also find that STDP increases the regularity of the interspike intervals (ISI) in spike trains of output neurons. However, the introduction of random refractory periods and fluctuating input currents can induce greater irregularity in the spike trains. Furthermore, when random inputs and STDP are combined, the correlation between the activity of different neurons increases. These findings suggest that the stochastic dynamics of STN neurons, in conjunction with STDP, could offer insights into the mechanisms underlying PD symptoms and their potential management.
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative disorder primarily characterized by the loss of dopaminergic neurons in the substantia nigra pars compacta, which leads to debilitating motor symptoms such as bradykinesia, rigidity, and tremors. Although various treatments, including pharmacological interventions and deep brain stimulation (DBS) of the subthalamic nucleus (STN), have shown efficacy in alleviating these symptoms, the mechanisms underlying these therapies remain poorly understood. Recent studies have highlighted the role of abnormal neuronal oscillations in the STN in contributing to motor dysfunction in PD, and suggested that synaptic plasticity-particularly spike-timing-dependent plasticity (STDP)-may play a crucial role in the modulation of these oscillations and the therapeutic effects of DBS [1, 2, 3].
STDP is a form of synaptic plasticity in which the strength of a synapse is adjusted based on the relative timing of pre- and postsynaptic spikes. This form of learning has been shown to be critical for maintaining neural network stability and promoting coordinated activity [4, 5]. In healthy brains, STDP is essential for processes like learning, memory, and sensory processing, but in the context of PD, its dysfunction may contribute to the altered neuronal activity observed in the STN [1, 6]. Moreover, the role of STDP in PD pathophysiology remains underexplored, especially with respect to how random factors-such as fluctuations in synaptic inputs or noise from sensory and thermal sources-affect neural firing patterns and network dynamics [7, 8]. This gap in understanding motivates our study, which aims to investigate how STDP interacts with stochastic factors in the dynamics of STN neurons. We develop a stochastic model that incorporates both random input currents and synaptic correlation, focusing on their combined effects on the cell membrane potential and neuronal spiking patterns in PD. While several studies have explored the effects of STDP and random inputs in neuronal networks, including modeling studies of the STN [9], as well as studies investigating the mechanistic features of networks with STDP and noise [10, 11], the specific interaction between these factors in the context of STN neurons remains underexplored. Additionally, related studies such as those by [12, 13] further elucidate the role of noise and plasticity in network dynamics. Burkitt et al. [14] also provide foundational insights into the general mechanisms of STDP in noisy environments, which are relevant for understanding the broader context of our study. Finally, additional work by Lindner [15] offers valuable perspectives on noise-driven network dynamics that complement our investigation. Specifically, we explore how STDP influences the regularity of interspike intervals (ISIs) and how random factors such as fluctuations in synaptic inputs or noise from sensory and thermal sources, neural firing patterns and network dynamics [16, 17]. This approach is distinct from prior studies, which typically examine STDP in more controlled, deterministic settings [18, 5].
Our model builds on well-established frameworks, including the Hodgkin-Huxley model for single-neuron dynamics [19] and integrate-and-fire (LIF) models for simplifying neuron spiking behavior [20], but uniquely integrates random input fluctuations and STDP to more accurately reflect the stochastic nature of biological neurons. Several studies have explored the interplay between noise and STDP in Hodgkin-Huxley (HH)-type neuron models [21] and in leaky integrate-and-fire (LIF) models (e.g., [22]). These studies provide important insights into how noise and plasticity together shape neuronal dynamics and learning processes. While previous work has focused on the effects of STDP in healthy networks, our study is one of the first to examine how STDP and random noise interact in the context of PD, a disease characterized by dysregulated neural circuits and irregular neuronal firing [23, 24]. Several studies have explored neural networks with STDP and noise in the context of PD. Chauhan et al. (2024) [25] investigated how synaptic weight and structural plasticity co-evolve, revealing how structural reorganization can enhance network synchrony and inform desynchronization strategies for PD. Madadi Asl et al. (2018) [26] showed that incorporating propagation delays and noise in STDP networks allows for bidirectional synapse preservation, offering insights into synaptic dynamics relevant to PD. These studies emphasize the importance of both excitatory and inhibitory inputs, which are crucial for accurately modeling the STN in PD. We build upon the deterministic STN model developed in [27], which was used to study the contributions of local cell and passing fiber activation to thalamic fidelity during DBS and lesioning. While their model considered deterministic dynamics, we introduce a stochastic term to capture the inherent random fluctuations in neuronal activity. These fluctuations, which arise from sensory noise, brainstem discharges, and thermal energy, are crucial for understanding the irregular firing patterns and synaptic plasticity observed in conditions like PD. By extending the model with this stochastic component, we investigate how random inputs influence the neuronal behavior and the efficacy of therapies like DBS.
The main contribution of this work is twofold. First, we propose a novel computational framework that combines STDP with random inputs to model the effects of noise and synaptic plasticity in PD. Second, we explore how these dynamics affect the regularity of neuronal firing, with potential implications for both PD pathophysiology and therapeutic interventions such as DBS. Our results suggest that while STDP can enhance the regularity of interspike intervals (ISIs), the presence of random input currents and refractory periods can introduce significant irregularity in spike trains, which may contribute to the motor symptoms observed in PD [5, 28]. These findings provide new insights into the neural mechanisms of PD and the potential role of STDP in managing the disease, with implications for future neuromorphic systems designed to mimic these processes [29, 30].
@@ 2 Model description
The phenomenology of STDP is typically described by a biphasic exponentially decaying function and order-dependent, with pre-post spiking inducing LTP and LTD [31, 32]. The instantaneous change in synaptic weight, $\Delta W$, is given by:
@@
where $\Delta W$ represents the change in synaptic weight, and $A_{+}$ and $A_{-}$ determine the maximum extent of synaptic modification. These maximum changes occur when the timing difference between presynaptic and postsynaptic spikes is close to zero. The parameters $\tau_{+}$ and $\tau_{-}$ define the time windows of presynaptic-to-postsynaptic interspike intervals that lead to synaptic strengthening or weakening [33]. The latency between presynaptic and postsynaptic spikes ($\Delta t$) is defined as:
@@
where $t_{\text{pre}}$ and $t_{\text{post}}$ represent the times of the presynaptic and postsynaptic spikes, respectively. When the postsynaptic neuron fires after the presynaptic neuron (i.e., $\Delta t<0$), this leads to a positive change in synaptic weight ($\Delta W>0$).
This model captures the key phenomenon that repeated presynaptic spikes occurring a few milliseconds before postsynaptic action potentials induce long-term potentiation (LTP) of the synapse, whereas presynaptic spikes occurring after postsynaptic spikes result in long-term depression (LTD) at the synapse.
@@ 2.1 Keeping track of pre- and postsynaptic spikes
To implement STDP while considering different synapses, we must track the times of both presynaptic and postsynaptic spikes throughout the simulation. To achieve this, we introduce two variables to monitor the spike times: $M(t)$ for the postsynaptic neuron and $P(t)$ for the presynaptic neuron. These variables are used to track the spikes over a longer timescale than the synaptic conductances. For each postsynaptic neuron, we define the following differential equation to track the postsynaptic spike history (see, e.g., [22]):
$\tau_{-}\frac{dM}{dt}=-M,$
where $\tau_{-}$ is the timescale over which the postsynaptic spike influence decays. Whenever the postsynaptic neuron spikes, we update $M(t)$ as follows:
@@
This allows $M(t)$ to track the number of postsynaptic spikes over the timescale $\tau_{-}$, and ensures that $M(t)$ is always negative, which is associated with the induction of LTD. Similarly, for each presynaptic neuron, we define the following differential equation:
$\tau_{+}\frac{dP}{dt}=-P,$
where $\tau_{+}$ is the timescale over which the presynaptic spike influence decays. When the presynaptic neuron spikes, we update $P(t)$ as follows:
@@
Here, $P(t)$ is always positive, corresponding to LTP. The variables $M(t)$ and $P(t)$ play similar roles to synaptic conductances $g_{i}(t)$, but instead of controlling the synaptic strength, they track the temporal relationship between presynaptic and postsynaptic spikes over much longer timescales. Importantly, $M(t)$ is negative and associated with LTD, while $P(t)$ is positive and associated with LTP, since they are updated by $A_{-}$ and $A_{+}$, respectively.
@@ 2.2 Implementation of STDP
The peak synaptic conductance of each synapse $i$, denoted by $\bar{g}_{i}$, will be adjusted based on the timing of presynaptic and postsynaptic spikes, using the variables $M(t)$ and $P(t)$. The conductance $\bar{g}_{i}$ can vary between $[0,\bar{g}_{\text{max}}]$ and is modified according to the relative timing of the spikes [22].
@@
    When the i-th presynaptic neuron fires a spike, the peak conductance is updated as follows:
@@
    Here, $M(t)$ tracks the time since the last postsynaptic spike and is always negative. Therefore, if the postsynaptic neuron spikes shortly before the presynaptic neuron, the peak conductance will decrease, as indicated by the negative value of $M(t)$.
@@
    When the postsynaptic neuron fires a spike, the peak conductance of each synapse is updated as follows:
@@
    Here, $P(t)$ tracks the time since the last spike of the i-th presynaptic neuron and is always positive.
    Thus, if the presynaptic neuron spikes before the postsynaptic neuron, the peak conductance increases, as indicated by the positive value of $P(t)$.
@@ 2.3 The leaky integrate-and-fire neuron connected with synapses that show STDP
We connect $N$ presynaptic neurons to a single postsynaptic neuron. We do not need to simulate the dynamics of each presynaptic neuron as we are only concerned about their spike times. So, we will generate $N$ Poisson-type spikes.
@@
where $I_{\text{app}}$ is the external input current, while $C_{m}$ is the membrane capacitance and $t\in[0,T]$ (some fixed time $T$). Additionally, in (11), $V_{\text{th}}$ denotes the membrane potential threshold to fire an action potential. In this model, we assume that a spike takes place whenever $V(t)$ crosses $V_{\text{th}}$ in the STN membrane potential. In that case, a spike is recorded and $V(t)$ resets to $V_{\text{reset}}$ value. Hence, the reset condition is summarized by $V(t)=V_{\text{reset}}$ if $V(t)\geq V_{\text{th}}$. The quantity $I_{\text{ahp}}$ represents the calcium-activated potassium current for the spike after hyperpolarization in STN.
@@
In our model (15)–(27), as we mentioned above, we use the simplest input spike train with Poisson process in which the stochastic process of interest provides a suitable approximation to stochastic neuronal firings [44]. The input spikes will be carried out by the quantity $\sum_{k}\delta(t-t_{k})$ in the equation (8) and each input spike arrives independently of the others, meaning that the timing of one spike does not influence the timing of subsequent spikes. The process will be described as follows:
@@
In this study, we model the external input to neurons using Poisson spike trains, which are commonly employed in computational neuroscience to simulate random background activity [45, 46]. Although real neurons show inter-spike intervals that deviate from the Poisson assumption [47, 48], Poisson processes remain a useful tool for isolating the effects of stochastic noise on STDP [49]. Poisson spike trains are simple and effective in providing random, uncorrelated inputs, which allows us to explore the basic dynamics of STDP and noise interactions in the context of PD without introducing additional complexities like input correlations or network-level dynamics.
In general, the information on stimulating activities in a neuron can be provided by the irregularity of spike trains. The time interval between adjacent spikes is called the ISI. The coefficient of variation (CV) of the ISI in a cell membrane potential with multiple inputs can bring useful information about the output of a decoded neuron. In what follows, we will demonstrate that when we increase the value of $\sigma_{\text{ref}}$, the irregularity of the spike trains increases (see also [50]). The term spike irregularity refers to the coefficient of variation of interspike intervals, which quantifies the variability in the timing of spikes rather than the variability of the spike count. The spike irregularity of spike trains can be described via the coefficient of variation of the inter-spike-interval (see, e.g., [51, 50]) as follows:
@@
Current Gating variables Gating variables Parameters
$I_{\text{L}}=g_{L}(v-E_{\text{L}})$ $g_{L}=2.25$ (nS)
$E_{\text{L}}=-60$ (mV)
$I_{\text{Na}}=g_{\text{Na}}m_{\infty}^{3}(V)h(V)(V-E_{\text{Na}})$ $m_{\infty}(V)=1/(1+\exp(-\frac{V+30}{15}))$ $h_{\infty}(V)=1/(1+\exp(-\frac{V+39}{3.1})$ $g_{\text{Na}}=37$
$\tau_{h}(V)=1+500/(1+\exp(-\frac{V+57}{-3})$ $E_{\text{Na}}=55$ (mV)
$I_{\text{K}}=g_{\text{K}}n^{4}(V)(V-E_{\text{K}})$ $n_{\infty}(V)=1/(1+\exp(-\frac{V+32}{8}))$ $g_{\text{K}}=45$ (nS)
$\tau_{n}(V)=1+100/(1+\exp(-\frac{V+80}{-26})$ $E_{\text{K}}=-80$ (mV)
$I_{\text{T}}=g_{\text{T}}a^{3}_{\infty}(V)b_{\infty}^{2}(r)r(V)(V-E_{\text{T}})$ $a_{\infty}(V)=1/(1+\exp(-\frac{V+63}{7.8}))$ $r_{\infty}(V)=1/(1+\exp(\frac{V+67}{2})$ $g_{\text{T}}=0.5$ (nS)
$b_{\infty}(V)=1/(1+\exp(-\frac{V-0.4}{0.1}))$ $\tau_{r}(V)=7.1+17.5/(1+\exp(-\frac{V+68}{-2.2})$ $E_{\text{T}}=0$ (mV)
$-1/(1+\exp(4))$
$I_{\text{Ca}}=g_{\text{Ca}}c^{2}(V)(V-E_{\text{Ca}})$ $c_{\infty}(V)=1/(1+\exp(-\frac{V+20}{8}))$ $g_{\text{Ca}}=2$ (nS)
$\tau_{c}(V)=1+10/(1+\exp(\frac{V+80}{26})$ $E_{\text{Ca}}=140$ (mV)
$I_{\text{ahp}}=g_{\text{ahp}}(V-E_{\text{ahp}})(\frac{\text{Ca}}{\text{Ca}+15})$ $g_{\text{ahp}}=20$ (nS)
$E_{\text{ahp}}=-80$ (mV)
$I_{\text{dbs}}=5+5\sin(2\pi t)$ (pA)